The equation y=5sin(3x-4t) where y is in millimeters, x is in meters and t is in seconds represents a wave motion.Determine the frequency
1
Expert's answer
2020-03-02T09:12:11-0500

The general form of wave equation


"y(x,t)=A\sin(kx-\omega t)."

We have


"y(x,t)=5\sin(3x-4 t)."

Hence, the angular frequency


"\omega=4\:\rm s^{-1}."

The frequency


"f=\omega/2\pi=4/2\pi=0.64\:\rm Hz."